# DART-Math: Difficulty-Aware Rejection Tuning for Mathematical Problem-Solving

### v1

@@ Abstract
Solving mathematical problems requires advanced reasoning abilities and presents notable challenges for large language models. Previous works usually synthesize data from proprietary models to augment existing datasets, followed by instruction tuning to achieve top-tier results. However, our analysis of these datasets reveals severe biases towards easy queries, with frequent failures to generate any correct response for the most challenging queries. Hypothesizing that difficult queries are crucial to learn complex reasoning, we propose Difficulty-Aware Rejection Tuning  (DART), a method that allocates difficult queries more trials during the synthesis phase, enabling more extensive training on difficult samples. Utilizing DART, we have created new datasets for mathematical problem-solving that focus more on difficult queries and are substantially smaller than previous ones. Remarkably, our synthesis process solely relies on a 7B-sized open-weight model, without reliance on the commonly used proprietary GPT-4. We fine-tune various base models on our datasets ranging from 7B to 70B in size, resulting in a series of strong models called DART-Math. In comprehensive in-domain and out-of-domain evaluation on 6 mathematical benchmarks, DART-Math outperforms vanilla rejection tuning significantly, being superior or comparable to previous arts, despite using much smaller datasets and no proprietary models. Furthermore, our results position our synthetic datasets as the most effective and cost-efficient publicly available resources for advancing mathematical problem-solving. [^1] Our datasets, models and code are publicly available at https://github.com/hkust-nlp/dart-math.
@@ 2.1 Background: Rejection Sampling and Rejection Tuning
We begin by formulating the data synthesis setting used for instruction tuning. For instruction tuning, the training dataset consists of $(x,y)$ pairs, where $x$ is the input query and $y$ is the response. The process of data synthesis involves generating new $(x,y)$ pairs to augment the original training dataset, thereby enhancing performance. For each input query $x_{i}$, it is typical to sample $M$ responses from advanced models such as GPT-4, forming the set $\{(x_{i},y_{i}^{(j)})\}_{j=1}^{M}$. In the context of mathematical problem-solving, a subsequent filtering step is often implemented to eliminate incorrect $y_{i}^{(j)}$. This elimination is based on whether the final answer in the synthetic response aligns with the ground-truth answer. [^2] Strictly speaking, final answer correctness does not necessarily imply intermediate reasoning correctness. We do not make further distinction across this paper which is not our focus. This is crucial as mathematical reasoning poses a significant challenge for current LLMs, and the generated $y_{i}^{(j)}$ may often be of poor quality. This method of response sampling is known as rejection sampling, and the subsequent fine-tuning process is referred to as rejection tuning, which is widely employed to enhance the mathematical problem-solving abilities of LLMs (Yuan et al., 2023; Yu et al., 2024; Singh et al., 2023; Xu et al., 2024). In addition to response synthesis, the queries are typically kept constant (Singh et al., 2023; Hosseini et al., 2024; Toshniwal et al., 2024) or altered in a controlled manner (Yu et al., 2024) to ensure that ground-truth answers are readily available, which facilitates the implementation of rejection sampling. While some studies also synthesize queries without utilizing rejection tuning (Li et al., 2024; Tang et al., 2024), our focus in this work is primarily on rejection tuning, a method prevalently used for advancing the mathematical skills of LLMs.
@@ 2.2 On the Imbalance of Rejection-Based Data Synthesis
3 DART – Difficulty-Aware Rejection Tuning
@@ 3.1 Open-Weight Models Are Able to Generate Good Responses
3.2 DARS – Difficulty-Aware Rejection Sampling
Motivated by the observation above, we aim to collect more responses for harder queries. Specifically, we introduce two strategies to increase the number of correct responses for difficult queries: (1) Uniform, which involves sampling responses for each query until each query accumulates $k_{u}$ correct responses, and $k_{u}$ is a preset hyperparameter determined by the desired size of the synthetic dataset; (2) Prop2Diff, where we continue sampling responses until the number of correct responses for each query is proportional to its difficulty score. The most challenging queries will receive $k_{p}$ responses and $k_{p}$ is a hyperparameter. This method introduces a deliberate bias in the opposite direction to vanilla rejection sampling, towards more difficult queries. Prop2Diff is inspired by previous works that demonstrate difficult queries can be more effective to enhance model capabilities (Sorscher et al., 2022; Liu et al., 2024b). Both the Uniform and Prop2Diff strategies prescribe a specific number of correct response for each query, determined by $k_{u}$ or $k_{p}$. Nevertheless, there are certain queries which we cannot sample out the designated number of correct responses even with extensive sampling efforts. To avoid endless running of the synthesis, we impose a cap on the maximum allowable number of raw samples per query as $n_{\max}$ – once this limit is reached for a particular query, we cease further sampling and retain any correct responses that have been gathered. The straightforward implementation of the Prop2Diff strategy risks generating no synthetic responses for easier queries if $k_{p}$ is set small. To mitigate this, we guarantee at least one synthetic response for each query when implementing Prop2Diff. While it might seem sufficient to rely on the original, real training dataset to ensure at least one human-annotated response per query, our findings highlight the importance of maintaining synthetic response coverage to learn to solve easy problems, as we will quantitatively shown in §4.3, partially because the human-annotated response is less detailed and not as beneficial as synthetic responses, demonstrated previously in Yu et al. (2024). For both Uniform and Prop2Diff strategies, we use the DeepSeekMath-7B-RL model to synthesize responses. We refer to the two sampling strategies as DARS-Uniform and DARS-Prop2Diff respectively. DARS-Prop2Diff requires assessing difficulties of queries, next we introduce an automatic approach to measure difficulties.
@@ 3.2 DARS – Difficulty-Aware Rejection Sampling
Previous studies have used proprietary models like ChatGPT to assess the difficulty or complexity of data samples (Lu et al., 2024; Liu et al., 2024b). In this work, we introduce a new metric, fail rate – the proportion of incorrect responses when sampling $n_{d}$ responses for a given query – as a proxy for difficulty. This metric aligns with the intuition that harder queries less frequently yield correct responses. We utilize DeepSeekMath-7B-RL as the sampling model to evaluate difficulty across all experiments in the paper. Varying this sampling model to align with the generative model may further enhance performance, which we leave as future work. Notably, one of the benefits of fail rate is that it allows to reuse the sampled responses during difficulty evaluation as synthetic responses for dataset construction. See implementation details in Appendix A.2.
@@ 4.1 General Setup
Below we summarize the key setup details, while we include more information in Appendix A.
@@
We perform standard instruction tuning on our synthetic datasets DART-Math-Uniform and DART-Math-Hard, based on several base models including Llama3-8B (Meta, 2024), Mistral-7B (Jiang et al., 2023), and Llama3-70B as representatives of general models, and DeepSeekMath-7B (Shao et al., 2024) as the representative of math-specialized models. For simplicity, we keep most hyperparameters the same across different models and datasets, and tune only several key hyperparameters like learning rate and number of epochs, as detailed in Appendix A.1.
@@ 4.2 Main Results
DART-Math achieves superior or comparable performance to previous best models. Specifically, when compared with MetaMath, DART-Math wins greatly in all cases. Additionally, DART-Math-DSMath-7B achieves the state-of-the-art results for models sized 7-8B on challenging benchmarks such as MATH, OlympiadBench-Math, and TheoremQA. On average, DART-Math-Mistral-7B (Prop2Diff) surpasses Mistral-7B-MMIQC by 4.6 absolute points, despite using only a quarter of its training sample size. Compared with concurrent work KPMath-Plus which relies on GPT-4 and has not released either the data or the model, our approach slightly underperforms on Mistral-7B for GSM8K and MATH. However, DART-Math excels against it on DeepSeekMath-7B by a significant margin, utilizing around one-third of its training data size. The Xwin-Math models perform well on the GSM8K benchmark but fall behind DART-Math (Prop2Diff) on other challenging benchmarks overall, particularly with a more pronounced gap on 70B models – although we note that their models are based on Llama2 which is not very fair to compare with. Importantly, we fully open-source our datasets and models, designating both DART-Math-Uniform and DART-Math-Hard as the best-performing and most cost-effective public instruction tuning datasets available for advancing mathematical problem-solving.
@@ 4.3 Analysis
DART generally needs more sampling trials to synthesize the same size of dataset compared to vanilla rejection tuning, as discussed in §3.3. It is important to underline that the synthesis cost, although initially higher, is a one-time expense. Once the dataset is synthesized, it can be used by the community and us to train numerous models, effectively amortizing the cost. To provide a quantitative understanding of the synthesis cost, we consider two main factors: $n_{\max}$, the maximum allowable raw samples for each query, and $r$, the ratio of queries that achieve the designated number of responses. If $n_{\max}$ is set too high, sampling may continue indefinitely for particularly difficult or noisy queries, resulting in a high synthesis cost. Conversely, a too small $n_{\max}$ may result in many queries not gathering the sufficient number of correct responses, leading to a lower $r$. Figure 4 (Right) illustrates the total number of raw samples required to synthesize 585K examples and the query achieving ratio $r$ as we increase $n_{\max}$. When $n_{\max}$ reaches 2048, over 90% of the queries can collect the designated number of responses under DARS-Uniform, with a corresponding total number of samples around 5 million. To reach 90% achieving ratio for DARS-Prop2Diff, $n_{\max}$ needs to be at least 8K, and the total number of raw samples exceeds 15 million. In our experiments, we achieved an over 95% ratio $r$, sampling approximately 150 million samples in total, which required about 5 days running inference of DeepSeekMath-7B-RL on 32 NVIDIA A100 GPUs. Besides that synthesis is a one-time cost, we would like to emphasize the number of samples is not a fair metric to compare synthesis cost between different works – our synthesis model of 7B size is relatively inexpensive and fast to run, compared to the much more costly and slower GPT-4 used in most previous studies. Moreover, achieving a query ratio as high as 95% may not be necessary to reach good performance. A slightly lower ratio of 85% or 90% might not significantly impact performance but could substantially reduce the synthesis cost. We plan to explore this balance further in future work.
@@ 5 Discussion
We utilize fail rate as the difficulty metric, yet it may be sub-optimal. Other metrics such as direct scoring (Liu et al., 2024b), Elo ratings, or the minimum pretraining compute to train a model that can always answer correctly (Burns et al., 2023) may be further explored. DART-Math is limited by natural language reasoning, while it is shown that generating and executing code helps solve mathematical problems significantly (Zhou et al., 2024; Yue et al., 2024; Gou et al., 2024; Liao et al., 2024; Toshniwal et al., 2024) – we think the bias in vanilla rejection sampling also exists for code generation, and DART could be integrated to potentially improve code generation as well.
Appendix A Experimental Setup
A.1 Training Setup

### v2

# for Mathematical Problem-Solving

@@ Abstract
Solving mathematical problems requires advanced reasoning abilities and presents notable challenges for large language models. Previous works usually synthesize data from proprietary models to augment existing datasets, followed by instruction tuning to achieve top-tier results. However, our analysis of these datasets reveals severe biases towards easy queries, with frequent failures to generate any correct response for the most challenging queries. Hypothesizing that difficult queries are crucial to learning complex reasoning, we propose Difficulty-Aware Rejection Tuning (DART), a method that allocates difficult queries more trials during the synthesis phase, enabling more extensive training on difficult samples. Utilizing DART, we have created new datasets for mathematical problem-solving that focus more on difficult queries and are substantially smaller than previous ones. Remarkably, our synthesis process solely relies on a 7B-sized open-weight model, without reliance on the commonly used proprietary GPT-4. We fine-tune various base models on our datasets ranging from 7B to 70B in size, resulting in a series of strong models called DART-Math. In comprehensive in-domain and out-of-domain evaluation on 6 mathematical benchmarks, DART-Math outperforms vanilla rejection tuning significantly, being superior or comparable to previous arts, despite using much smaller datasets and no proprietary models. Furthermore, our results position our synthetic datasets as the most effective and cost-efficient publicly available resources for advancing mathematical problem-solving. [^1] Our datasets, models and code are publicly available at https://github.com/hkust-nlp/dart-math.
@@ 2.1 Background: Rejection Sampling and Rejection Tuning
We begin by formulating the data synthesis setting used for instruction tuning. For instruction tuning, the training dataset consists of $(x,y)$ pairs, where $x$ is the input query and $y$ is the response. The process of data synthesis involves generating new $(x,y)$ pairs to augment the original training dataset, thereby enhancing performance. For each input query $x_{i}$, it is typical to sample $M$ responses from advanced models such as GPT-4, forming the set $\{(x_{i},y_{i}^{(j)})\}_{j=1}^{M}$. In the context of mathematical problem-solving, a subsequent filtering step is often implemented to eliminate incorrect $y_{i}^{(j)}$. This elimination is based on whether the final answer in the synthetic response aligns with the ground-truth answer. [^1] Strictly speaking, final answer correctness does not necessarily imply intermediate reasoning correctness. We do not make further distinction across this paper which is not our focus. This is crucial as mathematical reasoning poses a significant challenge for current LLMs, and the generated $y_{i}^{(j)}$ may often be of poor quality. This method of response sampling is known as rejection sampling, and the subsequent fine-tuning process is referred to as rejection tuning, which is widely employed to enhance the mathematical problem-solving abilities of LLMs (Zelikman et al., 2022; Yuan et al., 2023; Yu et al., 2024; Singh et al., 2023; Xu et al., 2024). In addition to response synthesis, the queries are typically kept constant (Singh et al., 2023; Hosseini et al., 2024; Toshniwal et al., 2024) or altered in a controlled manner (Yu et al., 2024) to ensure that ground-truth answers are readily available, which facilitates the implementation of rejection sampling. While some studies also synthesize queries without utilizing rejection tuning (Li et al., 2024; Tang et al., 2024), our focus in this work is primarily on rejection tuning, a method prevalently used for advancing the mathematical skills of LLMs.
@@ 2.2 On the Imbalance of Rejection-Based Data Synthesis
3 DART — Difficulty-Aware Rejection Tuning
@@ 3.1 Open-Weight Models Are Able to Generate Good Responses
3.2 DARS — Difficulty-Aware Rejection Sampling
Motivated by the observation above, we aim to collect more responses for harder queries. Specifically, we introduce two strategies to increase the number of correct responses for difficult queries: (1) Uniform, which involves sampling responses for each query until each query accumulates $k_{u}$ correct responses, and $k_{u}$ is a preset hyperparameter determined by the desired size of the synthetic dataset; (2) Prop2Diff, where we continue sampling responses until the number of correct responses for each query is (linearly) proportional to its difficulty score. The most challenging queries will receive $k_{p}$ responses and $k_{p}$ is a hyperparameter. This method introduces a deliberate bias in the opposite direction to vanilla rejection sampling, towards more difficult queries. Prop2Diff is inspired by previous works that demonstrate difficult queries can be more effective to enhance model capabilities (Sorscher et al., 2022; Liu et al., 2024b). Both the Uniform and Prop2Diff strategies prescribe a specific number of correct response for each query, determined by $k_{u}$ or $k_{p}$. Nevertheless, there are certain queries which we cannot sample out the designated number of correct responses even with extensive sampling efforts. To avoid endless running of the synthesis, we impose a cap on the maximum allowable number of raw samples per query as $n_{\max}$ — once this limit is reached for a particular query, we cease further sampling and retain any correct responses that have been gathered. The straightforward implementation of the Prop2Diff strategy risks generating no synthetic responses for easier queries if $k_{p}$ is set small. To mitigate this, we guarantee at least one synthetic response for each query when implementing Prop2Diff. While it might seem sufficient to rely on the original, real training dataset to ensure at least one human-annotated response per query, our findings highlight the importance of maintaining synthetic response coverage to learn to solve easy problems, as we will quantitatively shown in §4.3, partially because the human-annotated response is less detailed and not as beneficial as synthetic responses, demonstrated previously in Yu et al. (2024). For both Uniform and Prop2Diff strategies, we use the DeepSeekMath-7B-RL model to synthesize responses. We refer to the two sampling strategies as DARS-Uniform and DARS-Prop2Diff respectively. Though most previous methods are difficulty-agnostic, a few methods try assigning more budget to more complex questions to boost coverage, such as ToRA (Gou et al., 2024) and MARIO (Liao et al., 2024). However, ToRA/MARIO mainly focus on improving coverage without managing the distribution explicitly, leading to datasets that may still bias towards easy queries, while DARS explicitly controls the final distribution of the training dataset, completely eliminating the bias and also achieving higher coverage on the hardest queries. For more details about the comparison, we refer readers to Appendix A. As DARS-Prop2Diff requires assessing difficulties of queries, next we introduce an automatic approach to measure difficulties.
@@ 3.2 DARS — Difficulty-Aware Rejection Sampling
Previous studies have used proprietary models like ChatGPT to assess the difficulty or complexity of data samples (Lu et al., 2024; Liu et al., 2024b). In this work, we introduce a new metric, fail rate — the proportion of incorrect responses when sampling $n_{d}$ responses for a given query — as a proxy for difficulty. This metric aligns with the intuition that harder queries less frequently yield correct responses. We utilize DeepSeekMath-7B-RL as the sampling model to evaluate difficulty across all experiments in the paper. Varying this sampling model to align with the generative model may further enhance performance, which we leave as future work. Notably, one of the benefits of fail rate is that it allows to reuse the sampled responses during difficulty evaluation as synthetic responses for dataset construction. See implementation details in Appendix B.2.
@@ 4.1 General Setup
Below we summarize the key setup details, while we include more information in Appendix B.
@@
We perform standard instruction tuning on our synthetic datasets DART-Math-Uniform and DART-Math-Hard, based on several base models including Llama3-8B (Meta, 2024), Mistral-7B (Jiang et al., 2023), and Llama3-70B as representatives of general models, and DeepSeekMath-7B (Shao et al., 2024) as the representative of math-specialized models. For simplicity, we keep most hyperparameters the same across different models and datasets, and tune only several key hyperparameters like learning rate and number of epochs, as detailed in Appendix B.1.
@@ 4.2 Main Results
DART-Math achieves superior or comparable performance to previous best models. Specifically, when compared with MetaMath, DART-Math wins greatly in all cases. Additionally, DART-Math-DSMath-7B achieves the state-of-the-art results for models sized 7-8B on challenging benchmarks such as MATH, OlympiadBench-Math, and TheoremQA. On average, DART-Math-Mistral-7B (Prop2Diff) surpasses Mistral-7B-MMIQC by 4.6 absolute points, despite using only a quarter of its training sample size. Compared with concurrent work KPMath-Plus which relies on GPT-4 and has not released either the data or the model, our approach slightly underperforms on Mistral-7B for GSM8K and MATH. However, DART-Math excels against it on DeepSeekMath-7B by a significant margin, utilizing around one-third of its training data size. The Xwin-Math models perform well on the GSM8K benchmark but fall behind DART-Math (Prop2Diff) on other challenging benchmarks overall, particularly with a more pronounced gap on 70B models — although we note that their models are based on Llama2 which is not very fair to compare with. Importantly, we fully open-source our datasets and models, designating both DART-Math-Uniform and DART-Math-Hard as the best-performing and most cost-effective public instruction tuning datasets available for advancing mathematical problem-solving.
Additional results:
For additional results, such as domain-wise performance on MATH and comparison to RL, we refer readers to Appendix C.
@@ 4.3 Analysis
DART generally needs more sampling trials to synthesize the same size of dataset compared to vanilla rejection tuning, as discussed in §3.3. It is important to underline that the synthesis cost, although initially higher, is a one-time expense. Once the dataset is synthesized, it can be used by the community and us to train numerous models, effectively amortizing the cost. To provide a quantitative understanding of the synthesis cost, we consider two main factors: $n_{\max}$, the maximum allowable raw samples for each query, and $r$, the ratio of queries that achieve the designated number of responses. If $n_{\max}$ is set too high, sampling may continue indefinitely for particularly difficult or noisy queries, resulting in a high synthesis cost. Conversely, a too small $n_{\max}$ may result in many queries not gathering the sufficient number of correct responses, leading to a lower $r$. Figure 4 (Right) illustrates the total number of raw samples required to synthesize 585k examples and the query achieving ratio $r$ as we increase $n_{\max}$. When $n_{\max}$ reaches 2048, over 90% of the queries can collect the designated number of responses under DARS-Uniform, with a corresponding total number of samples around 5 million. To reach 90% achieving ratio for DARS-Prop2Diff, $n_{\max}$ needs to be at least 8K, and the total number of raw samples exceeds 15 million. In our experiments, we achieved an over 95% ratio $r$, sampling approximately 150 million samples in total, which required running inference of DeepSeekMath-7B-RL for about 160 NVIDIA A100 GPU days. Besides that synthesis is a one-time cost, we would like to emphasize the number of samples is not a fair metric to compare synthesis cost between different works — our synthesis model of 7B size is relatively inexpensive and fast to run, compared to the much more costly and slower GPT-4 used in most previous studies. Moreover, achieving a query ratio as high as 95% may not be necessary to reach good performance. A slightly lower ratio of 85% or 90% might not significantly impact performance but could substantially reduce the synthesis cost. We plan to explore this balance further in future work.
@@ 5 Discussion
We utilize fail rate as the difficulty metric, yet it may be sub-optimal. Other metrics such as direct scoring (Liu et al., 2024b), Elo ratings, or the minimum pretraining compute to train a model that can always answer correctly (Burns et al., 2023) may be further explored. DART-Math is limited by natural language reasoning, while it is shown that generating and executing code helps solve mathematical problems significantly (Zhou et al., 2024; Yue et al., 2024; Gou et al., 2024; Liao et al., 2024; Toshniwal et al., 2024) — we think the bias in vanilla rejection sampling also exists for code generation, and DART could be integrated to potentially improve code generation as well.
Acknowledgments
We thank Zhiyuan Zeng, Wei Xiong and Chenyang Zhao for helpful discussions. Yuxuan is partially supported by Tsinghua University Initiative Scientific Research Program (Student Academic Research Advancement Program).
Appendix A Comparison to Methods Based on Non-vanilla Rejection Sampling
Though both ToRA and MARIO have not released their datasets and focus on mathematical problem-solving using code in addition to natural language, which is out of our scope and thus not comparable, we try to implement the natural-language versions of their data synthesis strategies, which are comparable with DARS.
A.1 DARS produces distributions not biased towards easy queries
The most important difference between DARS and ToRA/MARIO is how responses are distributed across various queries — while we adjust the distribution either to be uniform or to favor more difficult queries, rather than merely improving coverage, ToRA/MARIO mainly focus on improving coverage without managing the distribution explicitly, leading to datasets that may still bias towards easy queries.
As shown in Table 3, though the absolute numbers of responses are not directly comparable between different methods, distribution-wise we can see that ToRA/MARIO still produce fewer responses for difficult problems than the easy ones. This especially contrasts with DART-Math-Hard, which produces, for example, 10x more responses for the MATH Level 5 queries than for the GSM8K queries.
As demonstrated in Figure 4 (Left), a high coverage rate (VRT+Cover) alone does not guarantee superior performance.
Synthetic Dataset Size RPQ in RPQ in Level-wise MATH MATH
(k) GSM8K 1 2 3 4 5 Coverage
ToRA 72 5.03 5.01 4.99 4.95 4.77 3.84 93.4%
MARIO 29 2.02 2.01 1.98 1.94 1.89 1.57 91.3%
DART-Math-Uniform 585 39.93 40.00 40.00 39.80 39.54 37.14 99.6%
DART-Math-Hard 590 8.49 14.28 33.52 54.94 79.59 107.06 99.6%
Table 3: Comparison between datasets synthesized by methods based on non-vanilla rejection sampling. “RPQ” means the average number of responses per query. The ToRA and MARIO datasets here are implemented by us according to their papers’ descriptions, since the official implementations have not been open-sourced.
A.2 DARS achieves high coverage even on the hardest queries
It is worth noting that a relatively high total coverage on MATH training set does not mean that the hard queries are well covered. For example, the MetaMathQA-MATH-AnsAug dataset achieves 82.8% of coverage on the MATH training set with evenly allocated budgets yet still admits missing a significant portion of hard queries and biasing towards easy queries, as analyzed in Figure 2.
In footnote 3 we show the coverage rate across all the difficulty levels by different methods. The ToRA-Corpus-16k statistics show that it only covers 68% of the Level 5 MATH queries while DART-Math datasets cover 99.6%.
MATH training set coverage Total Level 1 Level 2 Level 3 Level 4 Level 5
ToRA-Corpus-16k-MATH 83.1% 97.7% 91.6% 86.5% 81.3% 68.0%
MetaMath-MATH-AnsAug 82.8% 98.1% 93.6% 86.7% 76.6% 48.9%
VRT Baseline 84.9% 99.6% 98.2% 95.2% 89.8% 62.9%
DART-Math-* 99.6% 100.0% 100.0% 99.9% 99.7% 99.1%
Table 4: MATH training set coverage rates across all the difficulty levels of different synthetic datasets. The numbers of ToRA-Corpus-16k-MATH are from their OpenReview page [^3] https://openreview.net/forum?id=Ep0TtjVoap. The two DART-Math-* datasets have the same coverage because of the “Cover” operation, which tries to ensure there is at least one correct response for each query.
A.3 Details of Re-implementing Data Synthesis Strategies of ToRA and MARIO
Here we supplement more details on how we replicate the ToRA/MARIO synthesis pipeline to conduct the analysis present in the general author rebuttal. Below we show in the format as “ToRA/MARIO’s method -> how we adapt similar spirits for a simpler replication” step by step (we use CoT format rather than tool-integrated reasoning for a fairer comparison with our datasets):
ToRA:
  1. 1.
    Once for each problem in MATH&GSM8K with GPT-4, keeping the correct responses. -> We follow this with GPT-4o mini [^4] For GPT-4o mini, we use the version of gpt-4o-mini-2024-07-18 by default.
  2. 2.
    10 trials for each problem not correctly answered by greedy decoding with GPT-4 and keeping up to 4 correct responses per problem (to form ToRA-Corpus-16k). -> We follow this with GPT-4o mini.
  3. 3.
    Training CodeLlama models on ToRA-Corpus-16k to perform rejection sampling next. -> To avoid additional training for a fairer comparison, we use DeepSeekMath-7B-RL to replace the trained CodeLLama models here to align with DART-Math.
    1. (a)
      64 trials for each problem in MATH&GSK8K with CodeLlama, getting 233k distinct correct responses. -> We follow this with DeepSeekMath-7B-RL, getting 733k distinct correct responses.
    2. (b)
      Correcting wrong responses by greedy decoding from the correct preceding portions (costing no more than 64 trials for each problem) with CodeLLaMA-34B, getting 69k corrected responses. -> We simplify this by re-sampling another up to 64 trials per problem for all the incorrect responses, getting 225k correct samples.
    3. (c)
      Both ToRA and our adaptation: Randomly selecting up to 4 correct responses per problem from steps (a) and (b).
  4. 4.
    Merge ToRA-Corpus-16k and data from step 3 to form the final training dataset of 69k responses. -> We exactly follow this to form the final dataset of 72k responses.
MARIO:
  1. 1.
    Greedy decoding using GPT3.5 and GPT-4 each once for MATH&GSM8K, getting two responses for each query, only correct ones are kept -> We follow this but use GPT-4o mini to sample two responses for each query.
  2. 2.
    Sampling for 2 trials for each problem not correctly answered in step 1 using GPT-4, only correct ones are kept -> We follow this with GPT-4o mini.
  3. 3.
    Manually correcting responses for part of the remaining problems, then tuning Llemma-34B on it to obtain a synthesis agent for next steps -> this involves human annotation and is not comparable to our approach. For simplicity, we adopt DeepSeekMath-7B-RL as the synthesis agent to align with the DART-Math datasets.
  4. 4.
    Sampling with 100 trials and keeping up to 4 correct responses per problem for the remaining unanswered MATH queries, achieving 93.8% coverage on MATH -> we follow this and achieve 91.3% coverage on MATH.
  5. 5.
    Sampling with 1 trial for new problems introduced by MetaMath and keeping correct ones -> this step introduces new prompts and would only skew the distribution of responses, if any, towards easy queries. We remove this step for simplicity, which would not affect our conclusion.
Appendix B Experimental Setup
B.1 Training Setup